# Explainable Data-driven Modeling of Adsorption Energy in Heterogeneous Catalysis

### v1

@@ 2.4 Correlation from Symbolic Regression
Symbolic regression is a machine learning-based regression method aimed at identifying an underlying mathematical expression that best describes the relationship between input features and output using genetic programming algorithms. In our study, we employed the PySR library, which is an open-source tool designed to find an interpretable symbolic expression that optimizes a specified objective 36.
Constraints are imposed during symbolic regression to mitigate overfitting. Binary operators are restricted to addition, multiplication, subtraction, and division. We adopt MAE as the loss function, aligning with our approach to evaluate the performance of shallow machine learning models used in SHAP analysis. Furthermore, we set the model to an optimized value of 5000 iterations to prevent overfitting and enable turbo mode, which utilizes the Loop Vectorization method to accelerate the search evaluation among a pool of equations.
Due to the diverse range of adsorbates and catalysts, establishing a universal trend between input features and adsorption energy can be challenging. To overcome this obstacle, We employed symbolic regression on a truncated dataset composed of adsorbate-catalyst systems containing hydrogen atoms as adsorbates and metal catalysts, mirroring the setup in Trasatti’s research 34. Using the equation proposed by Trasatti, we computed the MAE value to be 1.005 for our dataset.
We apply the PySR model, which consistently reveals a relationship resembling that found in Trasatti’s research: the square of the catalyst electronegativity is directly proportional to the adsorption energy. Even after multiple training iterations, this relationship persists. The relationship obtained indicates that the square of the catalyst electronegativity is directly proportional to the adsorption energy. Notably, the MAE values obtained from PySR are consistently lower than those from the reference equation, even at lower complexities, and decrease further as complexity increases. However, higher complexities introduce numerous additional terms to minimize the MAE value, thereby detracting from a clearer understanding of the inherent relationship between features and the output label. Table 3 displays the resulting PySR equations. It is noteworthy that we consistently obtained correlations akin to Trasatti’s research for smaller values of complexity, regardless of the attempts made. We obtain $\Delta E_{\text{ads}}=kX_{\text{cat}}^{2}$, where the coefficient k varies, but the relationship obtained remains consistent. This indicates the reliability and reproducibility of the observed correlation, offering valuable insights into the underlying physics governing adsorption phenomena.
Table 3: Comparison of PySR equations and reference equation (eV)
PySR Equations MAE
$\Delta E_{\text{ads}}=0.0523X_{\text{cat}}^{2}$ $0.971\pm 0.003$
$\Delta E_{\text{ads}}=0.0914(\Delta H_{f}^{2}+X_{\text{cat}})$ $0.956\pm 0.019$
$\Delta E_{\text{ads}}=\frac{0.0901X_{\text{cat}}^{2}-\frac{0.0901(X_{\text{ads}}+0.1259)}{-\Delta H_{f}-2.5733}+\frac{0.0790}{-X_{\text{cat}}+\frac{2.1088}{X_{\text{cat}}-2.1088}}}{X_{\text{ads}}}$ $0.958\pm 0.023$
$\Delta E_{\text{ads}}=0.1029\Delta H_{f}(\Delta H_{f}+0.6525)+0.1029X_{\text{cat}}-\frac{0.0102}{X_{\text{cat}}-2.8382-\frac{0.0239}{X_{\text{ads}}}}+\frac{0.0026}{X_{\text{cat}}-2.8872}$ $0.945\pm 0.035$
Reference Equation
$\Delta E_{\text{ads}}=0.0104\Delta H_{f}+192.9660(0.01036(X_{\text{cat}}-X_{\text{ads}}))^{2}$ 1.005±$\pm 0.000$
@@ 3 Conclusion
Our study achieves explainability for ML models through post-hoc SHAP analysis and Symbolic Regression. This method highlights the importance and correlations of features. By merging ML techniques with XAI, we bridge the gap between data-driven methods and domain-specific knowledge, aiming to elucidate the intricate relationship between input structural features and adsorption energy. Key findings reveal that adsorbate properties exert a greater influence than catalyst properties, with a notable correlation between local electronegativity and catalyst electronegativity for our dataset. This correlation underscores the significant association between surface and bulk electronegativity, which is crucial for understanding surface phenomena in adsorption. Our findings suggest that the electronegativity of the catalyst bulk can serve as a proxy for surface electronegativity in practical applications. XAI techniques provide insights into the rationale behind ML model predictions by highlighting feature importance and the impact of feature values on adsorption energy values. This facilitates focused attention on relevant features for optimizing catalyst design. Overall, our study establishes a robust framework integrating ML techniques with XAI, leveraging extensive datasets like OC20 to enhance catalyst design through improved model explainability.
@@ 4.1 Dataset and Feature selection
The dataset utilized in this study is sourced from the Open Catalyst 2020 (OC20) dataset, which comprises 872,000 trajectories obtained from Density Functional Theory (DFT) calculation results, making it suitable for training machine learning (ML) models. From these trajectories, we extracted final frames to obtain relaxed structures of adsorbate-catalyst systems. The OC20 dataset encompasses catalyst systems featuring a diverse array of reactive elements, spanning nonmetals, alkali metals, alkaline earth metals, metalloids, transition metals, and post-transition metals. In contrast, the adsorbate systems consist of 82 molecules categorized into oxygen or hydrogen, C1 molecules, C2 molecules, and nitrogen-containing molecules.
For this study as seen in Table 4, we curated a dataset for post-hoc analysis by extracting a subset of the OC20 dataset, specifically focusing on systems containing oxygen (O), hydrogen (H), and C1 category adsorbates. This focus on small molecules minimizes the impact of interatomic interactions within the adsorbate molecules, simplifying comparisons with previous theory-based studies. Following this selection process, the dataset was narrowed down to a subset consisting of 2,886 systems. Additionally, a subset of 994 structures consisting hydrogen atom as adsorbate and metal catalysts was selected for Symbolic Regression analysis to derive mathematical equations representing the relationships between input features and adsorption energy.
Table 4: Datasets overview
Dataset Structures
OC20 Dataset 460,328
post-hoc XAI analysis 2886
Symbolic Regression 994
To extract features from this structure data, ASE and Pymatgen packages are utilized, enabling the extraction of relevant information from the relaxed structures of each trajectory. The adsorption energy is calculated as the difference between potential energy and reference energy in the OC20 dataset. Ultimately, the structured data, comprising 13 input features and adsorption energy as the output label, is fed into the ML models.
@@ 4.2 SHAP approach
We use SHAP library to perform post-hoc XAI analysis. SHAP (SHapley Additive exPlanations) is a technique based on game theory that helps explain how machine learning models make decisions. It connects fair credit distribution with simple explanations by using classic Shapley values. Shapely values are calculated using the Shapley regression values, a concept from cooperative game theory. Shapley regression values are feature importances for linear models in the presence of multicollinearity. This method requires retraining the model on all feature subsets $S\subseteq F$. Mathematically, the Shapley value for a feature $i$ is calculated as follows:
@@
4.3 Symbolic Regression
We use PySR (Symbolic Regression using Python) open-source library to perform symbolic regression on the dataset 37. The core of PySR model’s algorithm involves iteratively evolving populations of mathematical expressions through tournament selection 38, mutation, crossover, simplification, and optimization 39, 40. These populations evolve independently, with occasional migration of individuals between them. The evolve-simplify-optimize loop refines the discovered equations by simplifying them to equivalent forms and optimizing constants. This iterative process continues until satisfactory equations representing the relationships between input features and output labels are discovered.
In this framework, the inputs consist of the dataset containing input features and corresponding output label i.e. adsorption energy, while the output of the PySR model is a set of mathematical equations that represent the relationships between the input features and output label. These equations serve as interpretable models that capture the underlying patterns and dependencies within the data.

### v2

@@ 2.4 Correlation from Symbolic Regression
Symbolic regression is a machine learning-based regression method aimed at identifying an underlying mathematical expression that best describes the relationship between input features and output. In our study, we employed the PySR and SISSO++ libraries, which are open-source tools designed to find an interpretable symbolic expression that optimizes a specified objective 35, 36. In our study, we selected the top five important features with the highest Shapely values from the dataset and aimed to derive meaningful mathematical expressions related to adsorption energy. After removing highly correlated features and conducting post-hoc XAI analysis, the chosen five features were: the number of adsorbate atoms (N), adsorbate electronegativity ($\chi_{\text{ads}}$), coordination number of the adsorbing atom ($CN_{\text{adatom}}$), local electronegativity ($\chi_{\text{local}}$), and catalyst electronegativity ($\chi_{\text{cat}}$). These features serve as the input for deriving interpretable, physically meaningful equations using PySR and SISSO++.
PySR generates candidate equations using genetic programming and evaluates them with a user-defined loss function (e.g., RMSE, MAE)37. SISSO++ employs a compressed-sensing-based approach to generate candidate equations by minimizing a user-defined loss function, allowing for interpretable and accurate models 36. Both models, PySR and SISSO++, were configured with similar settings to generate equations at different complexity levels. SISSO++ outperformed PySR on both metrics. For more details on the algorithms, refer to the Methods section.
In this framework, the inputs consist of the dataset containing selected important input features and corresponding output label i.e. adsorption energy, while the output of the PySR and SISSO++ models is a set of mathematical equations that represent the relationships between the input features and output label. These equations serve as interpretable models that capture the underlying patterns and dependencies within the data.
Table 3: Comparison of PySR equations and SISSO++ equations
PySR equations
$E_{\text{ads}}=(-0.26338\cdot N)$
$E_{\text{ads}}=\frac{\chi_{\text{local}}-N}{\chi_{\text{ads}}}$
$E_{\text{ads}}=\left(\chi_{\text{local}}+(-0.49684\cdot N)\right)-0.78192$
$E_{\text{ads}}=\frac{\chi_{\text{ads}}-\left(N+\frac{CN_{\text{adatom}}}{\chi_{\text{local}}}\right)}{\chi_{\text{ads}}}$
$E_{\text{ads}}=((\chi_{\text{ads}}-(N+(CN_{\text{adatom}}/\chi_{\text{local}})))/(\chi_{\text{ads}}-0.18890163))$
SISSO++ equations
$E_{\text{ads}}=c_{0}+a_{0}\cdot\left(\frac{\chi_{\text{local}}-N}{\chi_{\text{ads}}^{6}}\right)$
$E_{\text{ads}}=c_{0}+a_{0}\cdot\left(CN_{\text{adatom}}-\chi_{\text{ads}}+\left|CN_{\text{cat}}-N\right|\right)+a_{1}\cdot\left(\frac{\chi_{\text{local}}-N}{\chi_{\text{ads}}^{6}}\right)$
$E_{\text{ads}}=c_{0}+a_{0}\cdot\cos(\exp(\chi_{\text{ads}}))+a_{1}\cdot\left(\left|CN_{\text{cat}}-N\right|+CN_{\text{adatom}}\right)+a_{2}\cdot\left(\frac{\chi_{\text{local}}-N}{\chi_{\text{ads}}^{3}}\right)$
$E_{\text{ads}}=c_{0}+a_{0}\cdot\cos(\chi_{\text{ads}}^{2})+a_{1}\cdot\left|\left(\chi_{\text{ads}}+\chi_{\text{local}}\right)-(\chi_{\text{ads}}\cdot\chi_{\text{local}})\right|+a_{2}\cdot\left(CN_{\text{adatom}}-\chi_{\text{ads}}+\left|CN_{\text{cat}}-N\right|\right)+a_{3}\cdot\left(\frac{\chi_{\text{local}}-N}{\chi_{\text{ads}}^{6}}\right)$
$E_{\text{ads}}=c_{0}+a_{0}\cdot\left|\chi_{\text{ads}}-\chi_{\text{local}}-\exp(-CN_{\text{adatom}})\right|+a_{1}\cdot\sin(\chi_{\text{local}})^{6}+a_{2}\cdot\cos(\chi_{\text{ads}}^{2})+a_{3}\cdot\left(\left|CN_{\text{cat}}-N\right|+CN_{\text{adatom}}\right)+a_{4}\cdot\left(\frac{N}{\chi_{\text{ads}}\cdot\chi_{\text{ads}}^{3}}\right)$
The equations generated by PySR and SISSO++, as shown in Table 3, consist of five equations for each method, presented in increasing levels of mathematical complexity. PySR equations are simpler, utilizing basic arithmetic operations, while SISSO++ produces more complex equations with higher-order terms that capture subtle, non-linear relationships in the data. PySR begins with a single input feature at complexity 1 and gradually includes more features as complexity increases, integrating all key features by complexity 5 (Table 3). In contrast, SISSO++ uses a broader range of parameters from the outset, incorporating three features at complexity 1 and utilizing all features by complexity 3. This leads to faster convergence rates and enhances its applicability in computational modeling. Another notable advantage of SISSO++ is its ability to consider feature units, ensuring physical relevance and interpretability of the resulting equations. For this study, feature units were specified during the SISSO++ setup, providing context that PySR lacks.
Although SISSO++ and PySR exhibit relatively high root mean square error (RMSE) values of 0.8 eV and 1.2 eV, respectively (details in the SI), SISSO++ successfully captures the correlation between input features and adsorption energy, aligning with the findings of Gao et al.32. Gao’s work shows a direct proportionality between adsorption energy and the effective coordination number, as seen in Equation 2. The first term contains $\psi$ which is electronic descriptor, $CN_{cat}$ in second term is effective coordination number and $\alpha$, $\lambda$ and $\theta$ are system-specific parameters. Similarly, SISSO++ identifies this proportionality from the second dimension onward, where $E_{\text{ads}}$ is directly proportional to $|CN_{cat}|$, the effective coordination number (as seen in table 3). These results suggest that at appropriate complexity levels, SISSO++ effectively captures critical feature interactions and underlying relationships, improving predictive accuracy while maintaining alignment with established literature.
$E_{\text{ads}}=0.1\cdot\alpha\cdot\psi+\lambda\cdot CN_{cat}+\theta$ (2)
In addition to capturing important feature expressions, the SISSO++ library offers a deterministic approach to model discovery, ensuring reproducible results—an advantage over the genetic programming methods like PySR which are stochastic in nature. SISSO++ also supports multi-task learning, allowing it to solve multiple related tasks simultaneously and leverage shared information to improve performance across all tasks.
@@ 3 Conclusion
Our study successfully integrates explainability into machine learning models through post-hoc SHAP analysis and symbolic regression, highlighting the importance and correlations of key features. By merging machine learning techniques with XAI, we bridge the gap between data-driven approaches and domain-specific knowledge, offering insights into the intricate relationships between input structural features and adsorption energy. One key finding is the direct proportionality between effective coordination number and adsorption energy in catalytic systems. This is corroborated by multiple analytical approaches: SHAP analysis identified effective coordination number as one of the top five influential features, while symbolic regression method, particularly SISSO++, expressed this direct relationship in its generated equations. Notably, these findings align with and further support the research of Gao et al., underscoring the robustness of this relationship across methodologies.
Additional insights from feature-feature correlations via SHAP analysis revealed a strong correlation between catalyst and local electronegativity, and how adsorbate electronegativity decreases with increase in number of adsorbate atoms for the dataset. The XAI techniques provide crucial insights into ML model predictions by highlighting feature importance and the impact of feature values on adsorption energy. This facilitates focused attention on relevant features, enabling symbolic regression to derive better relationships that can be leveraged for future catalyst optimization. In summary, our study establishes a robust framework that integrates machine learning techniques with XAI, using extensive datasets such as OC20 to improve catalyst design through enhanced model explainability.
@@ 4.1 Dataset and Feature selection
The dataset utilized in this study is sourced from the OC20 dataset, which comprises 872,000 trajectories obtained from DFT calculation results, making it suitable for training ML models. From these trajectories, we extracted final frames to obtain relaxed structures of adsorbate-catalyst systems. The OC20 dataset encompasses catalyst systems featuring a diverse array of reactive elements, spanning nonmetals, alkali metals, alkaline earth metals, metalloids, transition metals, and post-transition metals. In contrast, the adsorbate systems consist of 82 molecules categorized into oxygen or hydrogen, C1 molecules, C2 molecules, and nitrogen-containing molecules.
For this study, we curated a dataset for post-hoc analysis by extracting a subset of the OC20 dataset, specifically focusing on systems containing oxygen (O), hydrogen (H), and C1 category adsorbates. This focus on small molecules minimizes the impact of interatomic interactions within the adsorbate molecules, simplifying comparisons with previous theory-based studies. Following this selection process, the dataset was narrowed down to a subset consisting of 24503 systems. To extract features from this structure data, ASE and Pymatgen packages are utilized, enabling the extraction of relevant information from the relaxed structures of each trajectory. The adsorption energy is retrieved from the OC20 dataset. Ultimately, the structured data, comprising 19 input features and adsorption energy as the output label, is fed into the ML models.
@@ 4.2 SHAP approach
We use SHAP library to perform post-hoc XAI analysis. SHAP is a technique based on game theory that helps explain how machine learning models make decisions. It connects fair credit distribution with simple explanations by using classic Shapley values. Shapely values are calculated using the Shapley regression values, a concept from cooperative game theory. Shapley regression values are feature importances for linear models in the presence of multicollinearity. This method requires retraining the model on all feature subsets $S\subseteq F$. Mathematically, the Shapley value for a feature $i$ is calculated as follows:
@@
4.3 PySR
PySR (Symbolic Regression using Python) is an open-source library to perform symbolic regression on the dataset 37. The core of PySR model’s algorithm involves iteratively evolving populations of mathematical expressions through tournament selection 38, mutation, crossover, simplification, and optimization 39, 40. These populations evolve independently, with occasional migration of individuals between them. The evolve-simplify-optimize loop refines the discovered equations by simplifying them to equivalent forms and optimizing constants. This iterative process continues until satisfactory equations representing the relationships between input features and output labels are discovered.
@@ 4.3 PySR
To further refine these equations and mitigate overfitting, we imposed constraints during symbolic regression. We adopted RMSE as the loss function to enable direct comparison with the SISSO++ library at different complexity levels. Given the large dataset, we optimized the model with 50 iterations and enabled Turbo mode, which uses the Loop Vectorization method to accelerate the evaluation among a pool of equations. The max_size parameter was set to 7, limiting equation complexity to 7, while the max_depth was set to 5, restricting the maximum depth of the mathematical expressions or trees. We set lower values for max_depth and max_size to prioritize simplicity and interpretability, ensuring that the generated equations remain easy to understand and less prone to overfitting. Additionally, we enabled the denoise option, allowing the model to filter out noise in the input data and focus on the underlying relation rather than fitting to noise.
4.4 SISSO++
SISSO++ is an advanced feature selection and model construction algorithm that combines two key methods: Sure Independence Screening (SIS) and Subspace Optimization (SO). The SIS method, crucial for handling high-dimensional problems, begins with a large pool of potential features or descriptors. It calculates the correlation between each feature and the target variable (in this case, adsorption energy), then ranks and screens features based on these correlations. A subset of the most promising candidates is selected, and an independence check is performed to avoid redundancy and multicollinearity. This approach effectively reduces the dimensionality of the problem by retaining only the most relevant and independent features. Following the SIS step, the SO method comes into play. It performs a combinatorial optimization within the subspace of selected features to identify the best descriptor (or combination of descriptors) for the target property. SISSO++ applies this SIS-SO process iteratively, refining the selected features through multiple rungs.
The SISSO++ model settings used in this work are optimized to match the PySR model configuration for direct comparison. A descriptor dimensionality of 5 is chosen to capture sufficient information while controlling model complexity. A descriptor dimensionality of 5 (desc_dim = 5) is chosen to capture sufficient information while controlling model complexity. In each iteration, 3 features are selected (n_sis_select = 3) to ensure the most relevant variables are prioritized. The model progresses through 2 rungs (max_rung = 2), refining the selected features iteratively. To account for residual errors, 1 additional descriptor is added (n_residual = 1). The dataset used is specified by data.csv, and a 20% fraction of the data is reserved for testing purposes (leave_out_frac = 0.2) to ensure robust model validation similar to PySR model. Only the top-performing model is stored (n_models_store = 1), and the intercept is not fixed (fix_intercept = false) to allow for more flexible linear fitting. These settings are selected to achieve a balance between model performance and computational efficiency.
5 Data and Software Availability statement
[…]